Completely factor the expression.
x^2 + 10x + 16
1
Expert's answer
2013-02-08T05:27:32-0500
x^2 + 10x + 16
Applying Vieta's theorem we get that x1x2=16, x1+x2=-10
Therefore x1=-8, x2=-2
So given expression is factored in the following way:
x^2 + 10x + 16=(x+2)(x+8)